OpenStudy (anonymous):

Variable y varies directly with x2, and y = 196 when x = 7. Enter the constant of variation in the quadratic variation equation. y = x2

OpenStudy (anonymous):

Please help!!!!

OpenStudy (anonymous):

I'll give medal

jimthompson5910 (jim_thompson5910):

"y varies directly with x^2" means we have some equation in the form y = k*x^2 for some constant k They want the constant of variation, so they want to know the value of k.

jimthompson5910 (jim_thompson5910):

y = k*x^2 196 = k*(7)^2 ... plug in x = 7 and y = 196 solve for k

OpenStudy (anonymous):

But I don't understand how to get the awnser:/

OpenStudy (anonymous):

Please help my test is about to time out:( @jim_thompson5910

jimthompson5910 (jim_thompson5910):

I can't help with tests

OpenStudy (anonymous):

I know that will you jusy help me find the awnser? I want to learn how to do this

OpenStudy (anonymous):

@jim_thompson5910